OpenStudy (anonymous):

16/64 reduce to lowest terms.

OpenStudy (anonymous):

\[\frac{16}{64}=\frac{1}{4}\]

OpenStudy (anonymous):

\[\frac{16}{64} = \frac{2^4}{2^6} = \frac{1}{2^2} = \frac{1}{4} = 0.25\]

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

thanks.